OpenStudy (anonymous):

HELP WILL FAN AND MEDAL ON A TIME LIMIT THIS IS MY LAST CHANCE HHHEEELLLPPP i DON'T HAVE MUCH TIME The vertices of a right triangle are (–6, 4), (0, 0), and (x, 4). What is the value of x? I CAN'T GET IT

OpenStudy (anonymous):

If you put the vertices on a coordinate plane the only right angle possible is at (0,4) therefore x=0

OpenStudy (anonymous):

OK CAN YOU HELP WITH A FEW MORE

OpenStudy (anonymous):

I can try :)

OpenStudy (anonymous):

can you take a look at my question too?

OpenStudy (anonymous):

Which ordered pairs are solutions to the equation? y = x – 4 A. (10, 4) and (5, 1) B. (10, 6) and (5, 1) C. (10, 6) and (1, 5) D. (10, 5) and (5, 1) Thanks sssssssssssoooooooooo much and yes I will

OpenStudy (anonymous):

I am going to yours right now

OpenStudy (anonymous):

I don't know how to do that one I'm sorry :(

OpenStudy (mathmale):

To all of you: Please note that it is not helpful in the long run to do the work of others for thYem and to hand out free answers. Kindly help others find their own problem solutions.

OpenStudy (mathmale):

@tuxadog12 : which problem do you want to focus on?

OpenStudy (anonymous):

the second one

OpenStudy (anonymous):

y=x-4

OpenStudy (anonymous):

AloneS

OpenStudy (anonymous):

@AloneS

OpenStudy (anonymous):

help

OpenStudy (anonymous):

I have 10 min

OpenStudy (anonymous):

I can't fail

OpenStudy (mathmale):

Which ordered pairs are solutions to the equation? y = x – 4 You are given various points, and are asked which points lie on the graph of y = x - 4. If I gave you (2, 3), what would x be? what would y be? x= ? y= ?

OpenStudy (anonymous):

x=2 y=3

OpenStudy (mathmale):

Yes. Now, substitute x=2 and y = 3 into y=x-4. Is the equation then true or false?

OpenStudy (anonymous):

2=3-4

OpenStudy (mathmale):

true or false?

OpenStudy (anonymous):

true idk

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

@mathmale my time is almost up and I still have more

OpenStudy (anonymous):

@ash_heather1

OpenStudy (anonymous):

@AloneS

OpenStudy (anonymous):

@likeabossssssss

OpenStudy (anonymous):

help

OpenStudy (anonymous):

HHHHHHHEEEEEEEEEEILLLLLLLLLLLPPPPPPPPPP I AM ABOUT TO FAIL

OpenStudy (anonymous):

on which question do you need help with? you've posted a lot.

OpenStudy (anonymous):

Which ordered pairs are solutions to the equation? y = x – 4 A. (10, 4) and (5, 1) B. (10, 6) and (5, 1) C. (10, 6) and (1, 5) D. (10, 5) and (5, 1)

OpenStudy (anonymous):

HELP I HAVE 10 MIN AND MORE ?

OpenStudy (mathmale):

I hope you're not waiting for answers. My job is to help you solve these problems. I asked you a question earlier, but don't see that you've answered it. You are given y = x - 4. Is (8,4) a solution or not a solution? substitute 8 for x and 4 for y and decide whether the equation is now true or false.

OpenStudy (anonymous):

TRUE???? WAS MY ANSWERE

OpenStudy (anonymous):

@

OpenStudy (anonymous):

@AloneS I really need your help I am about to fail plz

OpenStudy (anonymous):

Which ordered pairs are solutions to the equation? y = x – 4 A. (10, 4) and (5, 1) B. (10, 6) and (5, 1) C. (10, 6) and (1, 5) D. (10, 5) and (5, 1)

OpenStudy (anonymous):

I think that it is C but IDK

alones (alones):

Okay first let me see if i'm doing this rightI'm solving for solve for x. y=x−4 x−4=y x−4+4=y+4 x=y+4 Eh let me graph it

OpenStudy (anonymous):

ok

alones (alones):

so a is way off on the graphing line

OpenStudy (anonymous):

yes

alones (alones):

Let's look at d

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

it is not D i think

alones (alones):

when you put this in a graph i mean i plugged the equation in..and now i 'm putting your choices in and most of them don't land on that points

alones (alones):

yeah it's not D as you can see

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Three vertices of square ABCD are A(–1, –5), B(–3, –1), and C(1, 1). What are the coordinates of the fourth vertex?

OpenStudy (anonymous):

most of the time I could answer this but my mind is blank

OpenStudy (anonymous):

help

alones (alones):

Hmmm aint suree /: